# A new framework for nonlinear Kalman filters

### v6

@@ Abstract
The Kalman filter (KF) is a state estimation algorithm that optimally combines system knowledge and measurements to minimize the mean squared error of the estimated states. While KF was initially designed for linear systems, numerous extensions of it, such as extended Kalman filter (EKF), unscented Kalman filter (UKF), cubature Kalman filter (CKF), etc., have been proposed for nonlinear systems. Although different types of nonlinear KFs have different pros and cons, they all use the same framework of linear KF. Yet, according to what we found in this paper, the framework often gives overconfident and less accurate state estimations when the measurement functions are nonlinear. Therefore, in this study, we designed a new framework for nonlinear KFs and showed theoretically and empirically that the new framework estimates the states and covariance matrix more accurately than the old one. The new framework was tested on four different nonlinear KFs and five different tasks, showcasing its ability to reduce the estimation errors by several orders of magnitude in low-measurement-noise conditions, with only about a 10–90% increase in computational time. To the best of our knowledge, all existing types of nonlinear KFs can benefit from the new framework, and the benefit will increase as the sensors become more and more accurate in the future. As an example, EKF, the simplest nonlinear KF that was previously believed to work poorly for strongly nonlinear systems, can now perform better than the iterated extended Kalman filter (IEKF) and provide fast and fairly accurate state estimations with the help of the proposed new framework. The codes are available at https://github.com/Shida-Jiang/A-new-framework-for-nonlinear-Kalman-filters.
{IEEEkeywords}
Cubature Kalman filters, extended Kalman filters, Kalman filters, nonlinear filters, state estimation, unscented Kalman filters.
@@ 1 Introduction
In control theory, the Kalman filter (KF) is a state estimation algorithm designed for linear systems. In such a system, a set of linear state transition functions describes how the states evolve, and a set of linear measurement functions describes the measured values of certain linear combinations of states. It is assumed that both the state transition and measurement functions contain some additive noise. The KF optimally combines the system knowledge and measurements to minimize the sum of the states’ variances when both process and measurement noise follow a normal (Gaussian) distribution. Since its birth in about 1960, KF has achieved massive success in navigation [1], motion control [2], statistics [3], signal processing [4], etc.
@@
Both EKF and EKF2 require the calculation of the Jacobian matrices of the state transition functions and measurement functions, which may not always be accessible in real-life systems. In 1997, Simon et al. introduced another extension of KF to nonlinear systems that requires no direct calculation of the Jacobian matrix and avoids the accuracy loss caused by linearization [7]. In later literature, this extension is usually referred to as the UKF [8]. UKF approximates the distribution of the states after nonlinear transformations by sampling several points (called sigma points) around the estimated mean of the states. These points go through the same transformation as the mean, and a more accurate estimation of the states’ distribution after the transformation can be obtained with the help of these extra sample points. In UKF, the sigma points are selected deterministically based on each state’s standard deviation. This technique guarantees that the sigma points are evenly distributed around the central point and can provide a more accurate state estimation. Similar ideas have also been used in several other nonlinear KFs, including cubature Kalman filter (CKF) [9], quadrature Kalman filter [10], square-root quadrature Kalman filter [11], etc. All these filters use deterministic ways to find several additional points around the central point (the estimated states from the previous iteration) and use those points to approximate the first two moments of the states after the nonlinear transformation. All these methods have some links with EKF2. As pointed out by Gustafsson et al. in [12], when the sigma points are infinitely close to the center, and the hyperparameters in the UKF are set to specific values, UKF will give the exact estimation as the EKF2. A disadvantage of these nonlinear KFs is that the number of sampling points is at least twice the number of states. Therefore, the method can be increasingly more complicated than EKF as the number of states increases.
@@
While different extensions of the KF use various ways to approximate the first two moments of the states after a nonlinear transformation, they all follow the same framework as in the linear KF. In such a framework, the state estimations in each iteration can be divided into two steps: Predict and Update. In the “predict” step, the state and covariance estimations are predicted using the state transition function and estimates from the previous iteration. Then, in the “update” step, the measurements correct the state predictions with a feedback gain (also known as the Kalman gain), which is designed to minimize the trace of the states’ covariance matrix after the update. After the states are updated, the state covariance matrix is also updated, ending the present iteration. Note that the covariance matrix will have the smallest possible trace after the update since that’s how the Kalman gain is selected.
@@
While such a framework can guarantee the optimality of the state estimation for linear systems, it often leads to overconfident covariance estimation and less accurate state estimation when applied to nonlinear systems. The primary reason for such a phenomenon is the assumption that the Kalman gain can minimize the trace of the state covariance matrix, which is generally not guaranteed for systems with nonlinear measurement functions. In fact, as we will show later, the “update” step can sometimes make the state estimation of the nonlinear system less accurate, so the assumption about the optimality of the “update” step can be extremely harmful, as it makes the algorithm become overconfident in an inaccurate estimation.
@@
Unfortunately, so far, there is no general solution that can effectively prevent the nonlinear KF from overconfidence. Although it is possible to mitigate the problem by approximating the measurement function at higher orders (e.g., using EKF2 instead of EKF), these algorithms take a significantly longer computational time, and the problem will still, in general, exist. The lack of theoretical analysis for nonlinear KFs makes finding a general solution even harder. For those few existing theoretical analyses [12, 13, 14], the results are based on the assumption that the higher-ordered terms in the Taylor expansions of the measurement functions can be ignored. This assumption may not be reasonable because the predicted states can be very far away from the updated states, and ignoring higher-ordered terms can cause significant errors in the ”update” step.
@@
  • To the best of our knowledge, we are the first to give theoretical proof that the conventional framework for nonlinear KFs, which has been used for around sixty years, has the problem of underestimating the trace of the state covariance matrix.
@@ 2.1 General formulation of nonlinear Kalman filters
where $x_{k}$ are the actual states at the time step $k$, $u_{k}$ are the inputs at the time step $k$, and $w_{k}$ are the process noises that follow the zero-mean multivariate Gaussian distribution with covariance $Q_{k}$. The estimations of the states made in the previous step are denoted as $\hat{x}_{k-1|k-1}$. Since the actual states are unknown, the KF assumes that the estimation errors $\hat{x}_{k-1|k-1}-x_{k-1}$ follow the zero-mean multivariate Gaussian distribution with covariance $P_{k-1|k-1}$. Whichever nonlinear KF is used, the first step of each iteration is called “predict”, which predicts the states using (1) and the estimations made in the previous time step. The predicted states are denoted as $\hat{x}_{k|k-1}$, and the errors $\hat{x}_{k|k-1}-x_{k|k-1}$ are assumed to follow the zero-mean multivariate Gaussian distribution with covariance $P_{k|k-1}$.
@@
where $z_{k}$ are the measurements at time step $k$, and $v_{k}$ are the measurement noises that follow the zero-mean multivariate Gaussian distribution with covariance $R_{k}$. Given (2) and the predicted states, the measurements can also be predicted using a particular method specific to the type of nonlinear KFs. The predicted measurements are denoted as $\hat{y}_{k|k-1}$, whose errors $\hat{y}_{k|k-1}-h(x_{k})$ are assumed to follow the zero-mean multivariate Gaussian distribution with covariance $P_{y}$. The actual cross-covariance matrix between the predicted states and the predicted measurements is denoted as $P_{xy}$. The differences between the predicted and actual measurements are called the measurement residuals $\tilde{y}_{k}=z_{k}-\hat{y}_{k|k-1}$. If we assume that the measurement noises are independent from the predicted states, then $\tilde{y}_{k}$ should have zero mean, and their covariance $S_{k}$ can be calculated by:
$S_{k}=P_{y}+R_{k}$ (3)
@@
where $\hat{x}_{k|k}$ are the updated states, and $K$ is the Kalman gain. With (3,4), the covariance of $\hat{x}_{k|k}$ can be calculated as:
@@
KFs select the Kalman gain to minimize the trace of $P_{k|k}$. By taking the derivative of (5), we notice that the trace is minimized when:
$K=K_{op}=P_{xy}S_{k}^{-1}$ (6)
@@
$P_{k|k}=P_{k|k-1}-K_{op}S_{k}K_{op}^{T}$ (7)
Equations (47) illustrate the key idea of KFs, which is to update the states in a way that minimizes the trace of the state covariance matrix after the update. However, it should be noted that these equations cannot be directly applied to systems with nonlinear measurement functions. The primary reason is that the covariance matrices $P_{y,k},S_{k}$, and $P_{xy,k}$ are all state-dependent for these nonlinear systems. Since the actual states are unknown, the actual values for these covariance matrices are also unknown. In fact, as we will show next, some over-simplifications regarding these matrices are the key factors that make the conventional framework for nonlinear KFs overconfident and inaccurate.
@@ 2.2 The conventional framework and its problem
While calculating $P_{y,k},S_{k}$, and $P_{xy,k}$ at the actual state is impossible, it is possible to approximate their values based on the predicted states. We denote these approximated covariance matrices as $P_{y,k|k-1},S_{k|k-1}$, and $P_{xy,k|k-1}$ since they are all estimated based on the prediction. With these approximations, the Kalman gain can be selected as:
$K=P_{xy,k|k-1}S_{k|k-1}^{-1}$ (8)
Although the detailed process of calculating $P_{y,k|k-1},S_{k|k-1}$, and $P_{xy,k|k-1}$ differs, all existing types of nonlinear KFs directly replace $S_{k}$ and $P_{xy,k}$ in (5) with the estimated values. Namely, the conventional framework estimates the updated state covariance matrix by:
$P_{k|k}$ $=P_{k|k-1}-KS_{k|k-1}K^{T}$ (9)
$=P_{k|k-1}-P_{xy,k|k-1}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}$
@@
On the other hand, the actual value of the updated state covariance matrix $P_{k|k,ac}$ can be calculated by substituting (8) into (5). Namely:
$P_{k|k,ac}=$ $P_{k|k-1}+P_{xy,k|k-1}S_{k|k-1}^{-1}S_{k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}$ (10)
$-P_{xy,k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}-P_{xy,k|k-1}S_{k|k-1}^{-1}P_{xy,k}^{T}$
While guaranteeing that $P_{k|k,ac}=P_{k|k}$ holds for all combinations of $P_{xy,k|k-1}$ and $S_{k|k-1}$ is impossible, we would at least hope that $P_{k|k}$ is an unbiased estimation of $P_{k|k,ac}$ (i.e., $\operatorname{\mathbb{E}}(P_{k|k})=P_{k|k,ac}$). Unfortunately, we have the following theorem.
Theorem 1
Define $g(A,B)=\text{tr}(P_{k|k,ac})|_{P_{xy,k}=A,S_{k}=B}$. Suppose that $\operatorname{\mathbb{E}}(P_{xy,k|k-1})=P_{xy,k}$, $\operatorname{\mathbb{E}}(S_{k|k-1})=S_{k}$, and $P_{xy,k|k-1}$ and $S_{k|k-1}$ are independent from each other. When $S_{k|k-1}-S_{k}$ is small, $P_{k|k}$ calculated from (9) has the following characteristic:
$\operatorname{\mathbb{E}}[\text{tr}(P_{k|k})]$ $=\operatorname{\mathbb{E}}[g(P_{xy,k|k-1},S_{k|k-1})]$ (11)
$<g(P_{xy,k},S_{k})=\text{tr}(P_{k|k,ac})$
Proof 2.2.
Denote $P_{xy,k|k-1}=P_{xy,k}+\Delta P$, $S_{k|k}=S_{k}+\Delta S$. According to the definition, $\operatorname{\mathbb{E}}(\Delta P)=O,\operatorname{\mathbb{E}}(\Delta S)=O$. Therefore, we have:
$g(P_{xy,k|k-1},S_{k|k-1})-g(P_{xy,k},S_{k|k-1})$ (12)
$=$ $\text{tr}[P_{xy,k|k-1}(S_{k}+\Delta S)^{-1}\Delta S(S_{k}+\Delta S)^{-1}P_{xy,k|k-1}^{T}]$
$-\text{tr}[\Delta PS_{k|k-1}^{-1}(P_{xy}+\Delta P)^{T}+(P_{xy}+\Delta P)S_{k|k-1}^{-1}\Delta P^{T}]$
Note that $(S_{k}+\Delta S)^{-1}$ can be expanded as the Neumann series. When $\Delta S$ is small, we have the following approximation:
$(S_{k}+\Delta S)^{-1}\Delta S(S_{k}+\Delta S)^{-1}$ (13)
$\approx$ $(S_{k}^{-1}-S_{k}^{-1}\Delta SS_{k}^{-1})\Delta S(S_{k}^{-1}-S_{k}^{-1}\Delta SS_{k}^{-1})$
$\approx$ $S_{k}^{-1}\Delta SS_{k}^{-1}-2S_{k}^{-1}\Delta SS_{k}^{-1}\Delta SS_{k}^{-1}$
Since $\Delta P$ and $\Delta S$ are independent from each other and have zero mean, we have:
$\operatorname{\mathbb{E}}[g(P_{xy,k|k-1},S_{k|k-1})-g(P_{xy,k},S_{k})]$ (14)
$\approx$ $\operatorname{\mathbb{E}}[\text{tr}(-2P_{xy,k|k-1}S_{k}^{-1}\Delta SS_{k}^{-1}\Delta SS_{k}^{-1}P_{xy,k|k-1}^{T}$
$-2\Delta PS_{k|k-1}^{-1}\Delta P^{T})]<0$
Note that the derivation of (14) utilizes the fact that both $S_{k}^{-1}\Delta SS_{k}^{-1}\Delta SS_{k}^{-1}$ and $S_{k|k-1}^{-1}$ are positive semi-definite and symmetrical.

### v10

@@ Abstract
The Kalman filter (KF) is a state estimation algorithm that optimally combines system knowledge and measurements to minimize the mean squared error of the estimated states. While KF was initially designed for linear systems, numerous extensions of it, such as extended Kalman filter (EKF), unscented Kalman filter (UKF), cubature Kalman filter (CKF), etc., have been proposed for nonlinear systems over the last sixty years. Although different types of nonlinear KFs have different pros and cons, they all use the same framework of linear KF. Yet, according to our theoretical and empirical analysis, the framework tends to give overconfident and less accurate state estimations when the measurement functions are nonlinear. Therefore, in this study, we designed a new framework that can be combined with any existing type of nonlinear KFs and showed theoretically and empirically that the new framework estimates the states and covariance more accurately than the old one. The new framework was tested on four different nonlinear KFs and five different tasks, showcasing its ability to reduce estimation errors by several orders of magnitude in low-measurement-noise conditions. The codes are available at https://github.com/Shida-Jiang/A-new-framework-for-nonlinear-Kalman-filters
keywords:
Kalman filtering; nonlinear observer and filter design; extended Kalman filter; unscented Kalman filter; estimation theory; parameter and state estimation
thanks: This paper was not presented at any IFAC meeting. Corresponding author Shida Jiang Tel. +1 510-944-9339.
@@ 1 Introduction
In control theory, the Kalman filter (KF) is a state estimation algorithm designed for linear systems. In such a system, a set of linear state transition functions describes how the states evolve, and a set of linear measurement functions describes the measured values of certain linear combinations of states. It is assumed that both the state transition and measurement functions contain some additive noise. The KF optimally combines the system knowledge and measurements to minimize the expected value of the mean-squared error of the state estimation. Additionally, when the process noise, measurement noise, and state estimation errors all follow a multivariate normal (Gaussian) distribution, the state estimation given by the Kalman filter is also the maximum likelihood estimation [1]. Due to these desirable features, KF is often referred to as the optimal linear state estimator. Since its birth in about 1960, KF has achieved massive success in navigation [2], motion control [3], statistics [4], signal processing [5], etc.
@@
Both EKF and EKF2 require the calculation of the Jacobian matrices of the state transition functions and measurement functions, which may not always be accessible in real-life systems. In 1997, Simon et al. introduced another extension of KF to nonlinear systems that requires no direct calculation of the Jacobian matrix and avoids the accuracy loss caused by linearization [8]. In later literature, this extension is usually referred to as the UKF [9]. UKF approximates the distribution of the states after nonlinear transformations by sampling several points (called sigma points) around the estimated mean of the states. These points go through the same transformation as the mean, and a more accurate estimation of the states’ distribution after the transformation can be obtained with the help of these extra sample points. In UKF, the sigma points are selected deterministically based on each state’s standard deviation. This technique guarantees that the sigma points are evenly distributed around the central point and can provide a more accurate state estimation. Similar ideas have also been used in several other nonlinear KFs, including cubature Kalman filter (CKF) [10], quadrature Kalman filter [11], square-root quadrature Kalman filter [12], etc. All these filters use deterministic ways to find several additional points around the central point (the estimated states from the previous iteration) and use those points to approximate the first two moments of the states after the nonlinear transformation. All these methods have some links with EKF2. As pointed out by Gustafsson et al. in [13], when the sigma points are infinitely close to the center, and the hyperparameters in the UKF are set to specific values, UKF will give the same state estimation as the EKF2, although the covariance estimation can still be a bit different. A disadvantage of these nonlinear KFs is that the number of sampling points is usually twice the number of states. Therefore, the method can be increasingly more complicated than EKF as the number of states increases.
@@
While different extensions of the KF use various ways to approximate the first two moments of the states after a nonlinear transformation, they all follow the same framework as in the linear KF. In this paper, this framework is referred to as “the conventional framework” or “the old framework”. In such a framework, the state estimations in each iteration can be divided into two steps: Predict and Update. In the “predict” step, the state and covariance estimations are predicted using the state transition function and estimates from the previous iteration. Then, in the “update” step, the measurements correct the state predictions with a feedback gain (also known as the Kalman gain), which is designed to minimize the trace of the states’ covariance matrix after the update. After the states are updated, the state covariance matrix is also updated, ending the present iteration. Note that the covariance matrix will have the smallest possible trace after the update since that’s how the Kalman gain is selected.
@@
While such a framework can guarantee the optimality of the state estimation for linear systems, it often leads to overconfident covariance estimation and less accurate state estimation when applied to nonlinear systems. The primary reason for such a phenomenon is the assumption that the Kalman gain minimizes the trace of the state covariance matrix, which is generally not guaranteed for systems with nonlinear measurement functions. In fact, as we will show later, the “update” step can sometimes make the state estimation of the nonlinear system less accurate, so the assumption about the optimality of the “update” step can be extremely harmful, as it makes the algorithm become overconfident in an inaccurate estimation.
@@
Unfortunately, so far, there is no general solution that can effectively prevent the nonlinear KF from overconfidence. Although it is possible to mitigate the problem by approximating the measurement function at higher orders (e.g., using EKF2 instead of EKF), these algorithms take a significantly longer computational time, and the problem will still, in general, exist. The lack of theoretical analysis for nonlinear KFs makes finding a general solution even harder. For those few existing theoretical analyses [13, 14, 15], the results are based on the assumption that the higher-ordered terms in the Taylor expansions of the measurement functions can be ignored. This assumption may not be reasonable because the predicted states can be very far away from the updated states, and ignoring higher-ordered terms can cause significant errors in the “update” step.
@@
  • To the best of our knowledge, we are the first to give a mathematical proof that the conventional framework for nonlinear KFs, which has been used for around sixty years, generally has the problem of underestimating the trace of the state covariance matrix.
@@ 2.1 General formulation of nonlinear Kalman filters
where $x_{k}$ are the actual states at the time step $k$, $u_{k}$ are the inputs at the time step $k$, and $w_{k}$ are the process noises that follow a zero-mean multivariate distribution with covariance $Q_{k}$. The estimations of the states made in the previous step are denoted as $\hat{x}_{k-1|k-1}$. Since the actual states are unknown, the KF assumes that the estimation errors $\hat{x}_{k-1|k-1}-x_{k-1}$ follow a zero-mean multivariate distribution with covariance $P_{k-1|k-1}$. Whichever nonlinear KF is used, the first step of each iteration is called “predict”, which predicts the states using (1) and the estimations made in the previous time step. The predicted states are denoted as $\hat{x}_{k|k-1}$, and the errors $\hat{x}_{k|k-1}-x_{k}$ are assumed to follow a zero-mean multivariate distribution with covariance $P_{k|k-1}$.
@@
where $z_{k}$ are the measurements at time step $k$, and $v_{k}$ are the measurement noises that follow a zero-mean multivariate distribution with covariance $R_{k}$. Given (2) and the predicted states, the measurements can also be predicted using a particular method specific to the type of nonlinear KFs. The predicted measurements are denoted as $\hat{y}_{k|k-1}$, whose errors $\hat{y}_{k|k-1}-h(x_{k})$ are assumed to follow a zero-mean multivariate distribution with covariance $P_{y,k}$. The actual cross-covariance matrix between the predicted states and the predicted measurements is denoted as $P_{xy,k}$. The differences between the predicted and actual measurements are called the measurement residuals, denoted as $\tilde{y}_{k}$. Namely:
$\tilde{y}_{k}=z_{k}-\hat{y}_{k|k-1}$ (3)
If we assume that the measurement noises are independent from the predicted states, then $\tilde{y}_{k}$ should have zero mean, and their covariance $S_{k}$ can be calculated by:
$S_{k}=P_{y,k}+R_{k}$ (4)
@@
where $\hat{x}_{k|k}$ are the updated states, and $K$ is the Kalman gain. With (35), the covariance of $\hat{x}_{k|k}$ can be calculated as [16]:
@@
KFs select the Kalman gain to minimize the trace of $P_{k|k}$. By taking the derivative of tr$(P_{k|k})$ in (6), we find that the trace is minimized when:
$K_{op}=P_{xy,k}S_{k}^{-1}$ (7)
@@
$P_{k|k,op}=P_{k|k-1}-K_{op}S_{k}K_{op}^{T}$ (8)
Equations (58) illustrate the key idea of KFs, which is to update the states in a way that minimizes the trace of the state covariance matrix after the update. However, it should be noted that these equations cannot be directly applied to systems with nonlinear measurement functions. The primary reason is that $P_{y,k},S_{k}$, and $P_{xy,k}$ can only be calculated if the actual state and the distribution of the states are known. Since neither of these holds in practice, the actual values for these covariance matrices are also unknown. In fact, as we will show next, some over-simplifications regarding these matrices are the key factors that can make the conventional framework for nonlinear KFs overconfident and inaccurate.
@@ 2.2 The conventional framework and its problem
While calculating $P_{y,k},S_{k}$, and $P_{xy,k}$ at the actual state is impossible, it is possible to approximate their values based on the predicted states. Different nonlinear filters (e.g. UKF, EKF2, CKF) take different approaches to approximating these covariance matrices. In the subsequent analysis, we denote these approximated covariance matrices as ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ since they are all estimated based on the prediction. Although the detailed process of calculating ${\color[rgb]{1,0,0}P_{y,k|k-1}}$, ${\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ differs, all existing types of nonlinear KFs directly replace $S_{k}$ and $P_{xy,k}$ in (7) and (8) with the approximated values. Namely, the Kalman gain is selected as:
${\color[rgb]{1,0,0}K_{est}}={\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ (9)
and the state covariance matrix is updated by:
${\color[rgb]{1,0,0}P_{k|k,est}}$ $=P_{k|k-1}-{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}S_{k|k-1}}{\color[rgb]{1,0,0}K_{est}^{T}}$ (10)
$=P_{k|k-1}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$
@@
On the other hand, the actual value of the updated state covariance matrix ${\color[rgb]{1,0,0}P_{k|k,ac}}$ given the selected Kalman gain in (9) can be calculated by substituting (9) into (6). Namely:
${\color[rgb]{1,0,0}P_{k|k,ac}}=P_{k|k-1}+{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}S_{k}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$ (11)
$\quad-P_{xy,k}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}P_{xy,k}^{T}$
In other words, ${\color[rgb]{1,0,0}P_{k|k,ac}}$ plays the role of $P_{k|k}$ in (6), and ${\color[rgb]{1,0,0}P_{k|k,est}}$ plays the role of $P_{k|k,op}$ in (8). Intuitively, tr$({\color[rgb]{1,0,0}P_{k|k,ac}})$ tends to be larger than tr$({\color[rgb]{1,0,0}P_{k|k,est}})$ since tr$(P_{k|k,op})=\min$tr$(P_{k|k})\leq$ tr$(P_{k|k})$. However, this relationship of size is not always preserved after the approximations. For example, according to (10) and (11), when $P_{xy,k}=S_{k}=1$ and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}={\color[rgb]{1,0,0}S_{k|k-1}}=0.9$, we have ${\color[rgb]{1,0,0}P_{k|k,est}}-{\color[rgb]{1,0,0}P_{k|k,ac}}=0.1>0$. Nevertheless, in most cases in practice [17, 18], we do find that tr$({\color[rgb]{1,0,0}P_{k|k,ac}})>$tr$({\color[rgb]{1,0,0}P_{k|k,est}})$, meaning that using (10) will often (although not always) underestimate the actual covariance.
In order to make this statement more rigorous in math, we consider the approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ as noisy versions of $P_{y,k},S_{k}$, and $P_{xy,k}$. Specifically, we assume that ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ is a random matrix, ${\color[rgb]{1,0,0}P_{y,k|k-1}}$ is a random positive semidefinite matrix, and ${\color[rgb]{1,0,0}S_{k|k-1}}$ is a random positive definite matrix. Note that we mark all the random covariance matrices in red in this section for clarity. Additionally, we assume that the “noises” that arise in the approximations are unbiased, so these random matrices should satisfy:
$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{y,k|k-1}}]=P_{y,k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S_{k|k-1}}]=S_{k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}]=P_{xy,k}\end{cases}$ (12)
These additional definitions and assumptions are made so that we can consider the “average” case and give a more mathematically rigorous definition of the overconfidence caused by (10). Namely, we have the following theorem.
Theorem 1.
Suppose that ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ and ${\color[rgb]{1,0,0}S_{k|k-1}}$ are random matrices that satisfy (12). ${\color[rgb]{1,0,0}S_{k|k-1}}$ is symmetric and positive definite. Then, $P_{k|k,est}$ calculated from (10) satisfies:
$\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{k|k,ac}}-{\color[rgb]{1,0,0}P_{k|k,est}}]\succeq 0$ (13)
where $\succeq$” represents positive semidefiniteness, and the equality only holds when ${\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ is constant.
{pf}
Let ${\color[rgb]{1,0,0}P}={\color[rgb]{1,0,0}P_{xy,k|k-1}}S_{k}^{-\frac{1}{2}}$, ${\color[rgb]{1,0,0}S}=S_{k}^{-\frac{1}{2}}{\color[rgb]{1,0,0}S_{k|k-1}}S_{k}^{-\frac{1}{2}}$. According to (12),
$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P}]=\bar{P}=P_{xy,k}S_{k}^{-\frac{1}{2}}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S}]=I\end{cases}$ (14)
According to (10, 11),
${\color[rgb]{1,0,0}P_{k|k,ac}}-{\color[rgb]{1,0,0}P_{k|k,est}}$ (15)
$=$ ${\color[rgb]{1,0,0}P}{\color[rgb]{1,0,0}S^{-2}P^{T}}-\bar{P}{\color[rgb]{1,0,0}S^{-1}P^{T}}-{\color[rgb]{1,0,0}PS^{-1}}\bar{P}^{T}+{\color[rgb]{1,0,0}PS^{-1}P^{T}}$
$=$ $({\color[rgb]{1,0,0}PS^{-1}}-\bar{P})({\color[rgb]{1,0,0}PS^{-1}}-\bar{P})^{T}+{\color[rgb]{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}$
For any fixed unit vector $\bm{v}$, ${\color[rgb]{1,0,0}P^{T}}\bm{v}$ is a random vector with a mean of $\bar{P}^{T}\bm{v}$. According to Lemma 4 in Appendix A,
$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}P}{\color[rgb]{1,0,0}S^{-1}}{\color[rgb]{1,0,0}P^{T}}\bm{v}]\geq\bm{v}^{T}\bar{P}\bar{P}^{T}\bm{v}.$ (16)
Moving the right hand side to the left, the following must hold true:
$\mathbb{E}[{\color[rgb]{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}]\succeq 0$ (17)
Therefore, the expectation of (15) is always p.s.d. due to (17) and the fact that the first term in (15) is p.s.d. The equality only holds when ${\color[rgb]{1,0,0}PS^{-1}}$ is constant, which is equivalent to ${\color[rgb]{1,0,0}K_{est}}={\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ being constant (i.e., free from approximation errors).